OpenStudy (anonymous):

Two objects A and B of the same mass collide elastically. Before collision object A is moving with a speed of 500m/s and object B is at rest. After the collision the two objects move along perpendicular paths, with the path of object A making an angle of 60 degrees from the orginal direction. After the collision, what is the speed of each object?

OpenStudy (anonymous):

draw a diagram.. and which conservation principle you think you can use here?

OpenStudy (anonymous):

 And I think that we can use the momentum equation, \[m _{A}v _{f} + m _{B}v _{f} = m _{A}v _{i} + m _{B}v _{i}\] this will help us in someway ! because initial velocity for B2 is 0 and then we also use: \[\frac{ 1 }{ 2 }m _{B}v _{f}^{2} + \frac{ 1 }{ 2 }m _{A}v _{f}^{2} = \frac{ 1 }{ 2 }m _{B}v _{i}^{2} + \frac{ 1 }{ 2 }m _{A}v _{i}^{2}\]